The mean IQ score for 1500 students is 100, with a standard deviation of 15. Assuming the scores have a normal curve,

how many have an IQ between 85 and 115?
1
Expert's answer
2013-02-25T11:51:17-0500
z(115) = (115-100)/15 = 1
z(85) = (85-100)/15 = -1
We need to consult Z-table to find P-value with current Z values
P(85< x < 115) = P(-1< z < 1) = 0.6827
# of 1500 that have IQ between 85 and 115 = 0.6827*1500 ~ 1024